OpenStudy (anonymous):

Write the equation of a circle with a center at (2.5, –3.5) and a diameter of 12. Answer (x + 2.5)2 + (y – 3.5)2 = 12 (x – 2.5)2 + (y + 3.5)2 = 36 (x + 2.5)2 + (y – 3.5)2 = 36 (x – 2.5)2 + (y + 3.5)2 = 144

OpenStudy (cwrw238):

general form is (x - a)^2 + (y - b)^2 = r^2 where (a,b) is the center and r = radius

OpenStudy (cwrw238):

in your case a = 2.5 and b = -3.5 and r , the radius, = half the diameter = 12/2 = 6

OpenStudy (cwrw238):

if you plug these values into the general equation you'll have it

OpenStudy (anonymous):

(x + 2.5)2 + (y – 3.5)2 = 12

OpenStudy (cwrw238):

no - look at the right hand side of the general equation its r^2 ( or r * r) if r = 6 what is r*r ?

OpenStudy (anonymous):

36! so then (x + 2.5)2 + (y – 3.5)2 = 36 ?

OpenStudy (cwrw238):

no its (x - 2.5)(x - (-3.5)) = 36 (x - 2.5)(x + 3.5) = 36 note the double negative giving +3.5

OpenStudy (cwrw238):

the second option is the correct one